A simple pendulum is found to vibrate 50 times within 200s. When 1.5m of its legth is reduced to a certain length, it vibrates 50 times in 175s. Find the original legth of the pendulum


1
Expert's answer
2021-02-10T11:01:47-0500

Let's first find the period of vibrations of first and second pendulum:


"T_1=\dfrac{t_1}{N}=\dfrac{200\ s}{50}=4\ s,""T_2=\dfrac{t_2}{N}=\dfrac{175\ s}{50}=3.5\ s."

From the definition of the period of pendulum, we have:


"T_1=2\pi\sqrt{\dfrac{L}{g}},""T_2=2\pi\sqrt{\dfrac{L-l}{g}}."

Let's divide the first expression by the second one:


"\dfrac{T_1}{T_2}=\sqrt{\dfrac{Lg}{g(L-l)}},""\dfrac{4}{3.5}=\sqrt{\dfrac{L}{(L-l)}},""L=1.31(L-l),""L=\dfrac{1.31l}{0.31}=\dfrac{1.31\cdot1.5\ m}{0.31}=6.34\ m."